# Adaptive Variation-Resilient Random Number Generator for Embedded Encryption

### v1

@@ IV.2 Compact Post-Processing: The Mini Trivium
IV.3 Experimental Implementation and Results
To test the overall proposed adaptive RNG system with the LFSR and the Mini-Trivium, the design was experimentally implemented using an FPGA board (NI myRIO-1900), as depicted in Fig. 6 (b). The schematic of Fig. 6 (a) depicts the complete system implementation, showing the connections between various components of the system. The combination of the linear feedback shift register (LFSR) along with the digital-to-analog converter (DAC) is employed as the entropy source that generates an analog stochastic signal ($V_{Stochastic}$), similar to the signal produced by the sMTJ and its associated circuitry presented in the previous section. Once generated, $V_{Stochastic}$ is then fed into both the LPF and and the comparator analog circuits within the adaptive digitizer. The LPF and the comparator of the adaptive digitizer were implemented using off-the-shelf discrete components on a breadboard, nonetheless, they can also be implemented using integrated components.
Figure 6: (a) Schematic diagram of system blocks and interconnections. (b) Experimental implementation of the adaptive RNG system.
@@ IV.3 Experimental Implementation and Results
A sample output from the adaptive digitizer is shown in Fig. 7 (a). As shown in the figure, the output looks like a non-ideal digital signal with finite slopes, and this is due to the limited slew rate (SR) of the used operational amplifier (OpAmp) for the comparator [37]. To obtain more ideal digital waveforms, a comparator with a higher SR can be employed or designed. This output from the adaptive digitizer - obtained using the breadboard-implemented analog circuit shown in Fig. 6 (b) - is then passed through the Mini Trivium cipher for post-processing. Fig. 7 (b) shows the generated random bit stream from the adaptive digitizer output of Fig. 7 (a) after post-processing. The compact Mini Trivium post-processing circuit was implemented using the FPGA unit. The FPGA unit was programmed and tested using a LABVIEW environment.
@@
As the results of Table 2 show, the LFSR on its own fails 12 out 16 tests consistently in all 10 experiments, confirming that the LFSR on its own is a low-quality PRNG. On the other hand, when the same random bit stream generated by the LFSR goes through the adaptive RNG system, namely the DAC, adaptive digitizer and the Mini Trivium cipher, then the resulting random bit stream shows excellent statistical randomness properties, consistently passing all 16 tests during all conducted experiments. Not only do the random bit streams from the LFSR-based adaptive RNG pass the tests, but they do so with excellent p-values consistently above 0.2. The obtained P-values are indeed comparable to the results obtained for the sMTJ-based adaptive TRNG system at nominal conditions, as shown quantitatively through both Table 2 and the visual illustration of Fig. 8. The experimental results confirm that the LFSR-based adaptive RNG implementation using off-the-shelf components results in the generation of high quality random numbers with properties suitable for embedded encryption.
@@ V Conclusion
In this work, we demonstrated an adaptive RNG for variation-resilient extraction of random numbers from entropy sources based on emerging device technologies. Two demonstrations were presented, one employing an sMTJ-based entropy source, which was comprehensively evaluated using MATLAB + SPICE simulations, and the other employing an LFSR-based pseudo random entropy source that emulates the true-random source, which was experimentally implemented using an FPGA.
A key component of the adaptive RNG system was its adaptive digitizer, which employed a low-pass filter to generate an adaptive reference voltage that tracks any deviation or drift in the generated stochastic signal from the entropy source, enabling variation-resilience and allowing extraction of unbiased random bit streams with high entropy. Statistical properties of generated bit streams by the adaptive RNGs were tested using the NIST test suite, demonstrating encryption-grade randomness consistently. The presented adaptive RNG has great potential for use in embedded encryption within IoT devices, making data encryption accessible locally.

### v2

@@ IV.2 Compact Post-Processing: The Mini Trivium
IV.3 Experimental Implementation and Results
To test the overall proposed adaptive RNG system with the LFSR and the Mini-Trivium, the design was experimentally implemented using discrete electronic components and an FPGA board (NI myRIO-1900), as depicted in Fig. 6 (b). The schematic of Fig. 6 (a) depicts the complete system implementation, showing the connections between various components of the system. The combination of the linear feedback shift register (LFSR) along with the digital-to-analog converter (DAC) is employed as the entropy source that generates an analog stochastic signal ($V_{Stochastic}$), similar to the signal produced by the sMTJ and its associated circuitry presented in the previous section. Once generated, $V_{Stochastic}$ is then fed into both the LPF and and the comparator analog circuits within the adaptive digitizer. The LPF and the comparator of the adaptive digitizer were implemented using off-the-shelf discrete components on a breadboard, nonetheless, they can also be implemented using integrated components.
A sample output from the adaptive digitizer is shown in Fig. 7 (a). As shown in the figure, the output looks like a non-ideal digital signal with finite slopes, and this is due to the limited slew rate (SR) of the used operational amplifier (OpAmp) for the comparator [43]. To obtain more ideal digital waveforms, a comparator with a higher SR can be employed or designed. This output from the adaptive digitizer - obtained using the breadboard-implemented analog circuit shown in Fig. 6 (b) - is then passed through the Mini Trivium cipher for post-processing. Fig. 7 (b) shows the random bit stream generated from the output of the adaptive digitizer that was shown in Fig. 7 (a) after post-processing. The compact Mini Trivium post-processing circuit was implemented using the FPGA unit. The FPGA unit was programmed and tested using a LABVIEW environment.
@@ IV.3 Experimental Implementation and Results
Table 3: Performance Comparison of State-of-The-Art TRNGs
Reference and Year [22] [41] [11] [36] [9] [30] [57] [37] This
Parameter 2022 2023 2017 2019 2022 2023 2024 2025 Work
Entropy Source Technology CMOS CMOS CMOS RRAM RRAM MTJ MTJ sMTJ sMTJ
Entropy Source Mechanism Oscillator Jitter Oscillator Jitter Metastability and Jitter Random Telegraph Noise Intra- Device Switching Static Stochastic Switching Stochastic Initialization Switching Static Stochastic Switching Dynamic Stochastic Switching
Operating Voltage ($V_{DD}$) $1.2$ V $1.1$ V $1.2$ V $0.7$ V $5.0$ V
Switching Voltage ($V_{SET}$) $1.0$ V $2.0$ V
Throughput (Mbps) 52 53 3000 40 303 5–182
NIST Tests Pass Rate 15/15 15/15 15/15 15/15 15/15 14/14 16/16 15/15 16/16
Variation Resilient Yes Yes Yes Yes No Yes No Yes Yes
Post-Processing Included No Yes No Yes No No Yes No Yes
Off-Shelf Implementation Yes
Entropy Source Generality Yes
CMOS: Complementary Metal Oxide Semiconductor — RRAM: Resistive Random Access Memory — MTJ: Magnetic Tunnel Junction
Based on reported results in this paper but in principle can go to much higher throughput (up to Gbps) with faster stochastic MTJs [34, 33].
As the results of Table 2 show, the LFSR on its own fails 12 out 16 tests consistently in all 10 experiments, confirming that the LFSR on its own is a low-quality PRNG. On the other hand, when the same random bit stream generated by the LFSR goes through the adaptive RNG system (namely the combination of the DAC, adaptive digitizer and the Mini Trivium cipher) then the resulting random bit stream shows excellent statistical randomness properties, consistently passing all 16 tests during all conducted experiments. Not only do the random bit streams from the LFSR-based adaptive RNG pass the tests, but they do so with excellent P-values consistently above 0.2. The obtained P-values are indeed comparable to the results obtained for the sMTJ-based adaptive TRNG system at nominal conditions, as shown quantitatively through both Table 2 and the visual illustration of Fig. 8. The experimental results confirm that the LFSR-based adaptive RNG implementation using off-the-shelf components results in the generation of high quality random numbers with properties suitable for embedded encryption.
Finally, to benchmark our presented adaptive RNG with other implementations, a performance comparison of our adaptive RNG with the existing state-of-the art TRNG designs based on CMOS, RRAM and MTJ technologies is presented in Table 3. The presented adaptive RNG offers a wide - relatively high - throughput range of 5 Mbps to 182 Mbps ($\tau_{c}$ of $\mus - 2.5 μ\mus$ respectively), and a consistent 100% NIST test pass rate. It is important to mention here, that our approach does not need voltage pulsing as required in other designs, and also our design is implementable using off-shelf component implementations. It is important to also highlight the generality of our adaptive RNG concept beyond MTJs only, as it can be implemented using various entropy sources (e.g. CMOS, RRAM, MTJ), as demonstrated through our LFSR implementation in the FPGA experiments.
@@ V Conclusion
In this work, we demonstrated an adaptive RNG for variation-resilient extraction of random numbers from entropy sources based on emerging device technologies. Two demonstrations were presented, one simulation-based and one experimental, with the first employing an sMTJ-based entropy source, which was comprehensively evaluated using MATLAB + SPICE simulations, and the other employing an LFSR-based pseudo random entropy source that emulates the true-random source, which was experimentally implemented using discrete electronic components and an FPGA.
A key component of the adaptive RNG system was its adaptive digitizer, which employed a low-pass filter to generate an adaptive dynamically changing short-term-average reference voltage that tracks any deviation or drift in the generated stochastic signal from the entropy source, enabling variation-resilience and allowing extraction of unbiased random bit streams with high entropy. The presented adaptive RNG consistently passed all NIST SP 800-22 tests, and demonstrated a throughput in the range 5-182 Mbps, while maintaining high randomness qualities. Although the NIST test suite is the commonly used framework for evaluating randomness qualities of random bit streams generated using TRNGs, the presented analysis can be further expanded in the future to include other methods as well, such as the Dieharder [42], TestU01 [38, 28, 27] and PractRand [52], which can be useful for more comprehensive assessment of the randomness qualities of TRNGs for mission critical and highly sensitive encryption applications. The presented adaptive RNG has great potential for use in embedded encryption within IoT devices, making data encryption accessible locally.